Decomposition of h2o2 is followed by titration with kmno4. 5ml of reaction mixture needs 46ml of solution after 5min same vol meeds 37 ml of solution what fraction of solution decompose in 5min
 
 

Grade:12

1 Answers

Vikas TU
3 years ago
Dear student 
t = 2.303/k logV0/V1 
= 1/k ln v0/v1 
= ln5 / 4.5 * 10^-2 min 
Good Luck 
Cheers …...............................